OpenStudy (anonymous):

Is g(-x)=6x-x^3 odd, even, or neither?

OpenStudy (anonymous):

odd

OpenStudy (anonymous):

since the exponents are all odd

OpenStudy (anonymous):

Oh o.k. Thank you satellite for explaining :)

OpenStudy (anonymous):

\[g(x)=6x-x^3\] \[g(-x)=6(-x)-(-x)^3=-6x+x^3=-(6x-x^3)=-g(x)\] and if \[g(-x)=-g(x)\] then g is "odd"